OpenStudy (anonymous):

A table is on sale for 73%off. The sale price is $226.80 What is the regular price?

OpenStudy (anonymous):

discount is 73 % so one has to pay 27 % of the actual regular price. which means 27% of x = 226.80 => 27 *x / 100 = 226.80 => x = 22680/27 = 840 USD

OpenStudy (lgbasallote):

0.27x = 226.80 x = 226.80/0.27 x = 840

OpenStudy (anonymous):

yeh i totally got it wrong! vikash and igbasallote are correct. gd job guys